A coil ACD of N turns and radius R carries a current I Amp and is placed on a horizontal table.K is a very small horizontal conducting ring of radius r placed at a distance  Y  frm the centre of the coil vertically above the coil ACD .Find an  expression for the EMF estabilished when the ring K is allowed to fall freely.Express the emf in terms of instantaneous speed v and height Y.


Grade:12th Pass

1 Answers

ROSHAN MUJEEB
3 years ago

Solution :
B=μ0IR2N2[R2+y2]3/2B=μ0IR2N2[R2+y2]3/2
ε=−dϕdt=−AdBdt=−πr2dBdtε=-dϕdt=-AdBdt=-πr2dBdt...(1)
heredBdt=−μ0IR2N2×321[R2+y2]5/2×2y×dydtdBdt=-μ0IR2N2×321[R2+y2]5/2×2y×dydt
=−3μ0IR2N2yv[R2+y2]5/2=-3μ0IR2N2yv[R2+y2]5/2put this value in eq. (i) and get Ans.